OpenStudy (anonymous):

Solve the following proportion: x/9 = 48/36

OpenStudy (anonymous):

Let's solve your equation step-by-step. x9=4836 Step 1: Cross-multiply. x9=4836 x*(36)=(48)*(9) 36x=432 Step 2: Divide both sides by 36. 36x36=43236 x=12 Answer: x=12

OpenStudy (anonymous):

how are u so smart

OpenStudy (anonymous):

and thanks

OpenStudy (anonymous):

your welcome